OpenStudy (anonymous):

Solve the equation for y. 7x – 5y = 10

OpenStudy (alexandervonhumboldt2):

-5y=10-7x y=-(10-7x)/5 y=-2+7/5x y=7/5x-2

OpenStudy (alexandervonhumboldt2):

@chosenmatt